Find the 95% confidence interval for the standard deviation of the lengths of pipes if a sample of 11 pipes has a standard deviation of 14.6 inches.
Found 2 solutions by rothauserc, jim_thompson5910:
Answer by rothauserc(4718)

We use a chi-squared distribution which gives us a confidence interval that is not symmetric about the sample standard deviation
:
We assume that the population of pipes is normally distributed, then
:
chi-squared(1 - a/2) < or = (n-1)s^2/(std. dev.)^2 < or = chi-squared(a/2)
:
where a is the alpha, s is the sample standard deviation, and std. dev. is the population standard deviation
:
a = (1 - 95/100) = 0.05
:
for this problem a = 0.05, degrees of freedom = 11 - 1 = 10, then
:
chi-squared(0.975) = 20.5 and chi-squared(0.025) = 3.25
:
we evaluate (n-1)s^2 / chi-squared for 20.5 and 3.25
:
(11-1) * 14.6^2 / 20.5 = 103.98
(11-1) * 14.6^2 / 3.25 = 655.88
:
we take the square root of 103.98 and 655.88
:
square root(103.98) = 10.197 approx 10.2
square root(655.88) = 25.610 approx 25.6
:
*******************************************
the 95% confidence interval is (10.2, 25.6)
*******************************************
:

Answer by jim_thompson5910(35256)

Given Info:
Confidence Level = C = 95% = 0.95
sample size = n = 11
sample standard deviation = s = 14.6
---------------------------------------------------------------------------
Based on the sample size, the degrees of freedom (df) is
df = n-1 = 11-1 = 10

Let
L = left chi-square critical value
R = right chi-square critical value

We need to find the L & R values that make true. The symbol is the Chi-Square symbol

Using a table or calculator those approximate values are
L = 3.247
R = 20.483
For this problem, I used a table. Specifically I used this table. Look at the df = 10 row. Highlight the columns that have 0.975 and 0.025 up top. These values are areas to the right of the values marked in the table. So for example, . The middle area has 95% area. The two tails combine to 5% so each tail has area of 0.05/2 = 0.025
---------------------------------------------------------------------------
Now we can compute the confidence interval








---------------------------------------------------------------------------
The 95% confidence interval for the standard deviation is approximately (10.2, 25.6)